$SPAD/input schaum33.input

Timothy Daly December 3, 2016

Contents

2

1 [?]:14.636 ∫ csch ax dx

∫ csch ax = \frac{1}{a} ln tanh \frac{ax}{2}

— * —

)set break resume
)spool schaum33.output
)set message test on
)set message auto off
)clear all
--S 1 of 49
aa:=integrate(csch(a*x),x)
--R
--R
--R        - log(sinh(a x) + cosh(a x) + 1) + log(sinh(a x) + cosh(a x) - 1)
--R   (1)  -----------------------------------------------------------------
--R                                        a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 2 of 49
bb:=1/a*log(tanh((a*x)/2))
--R
--R                 a x
--R        log(tanh(---))
--R                  2
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 3 of 49
cc:=aa-bb
--R
--R   (3)
--R                  a x
--R       - log(tanh(---)) - log(sinh(a x) + cosh(a x) + 1)
--R                   2
--R     +
--R       log(sinh(a x) + cosh(a x) - 1)
--R  /
--R     a
--R                                                    Type: Expression(Integer)
--E
--S 4 of 49      14:636 Schaums and Axiom agree

dd:=complexNormalize cc
--R
--R   (4)  0
--R                                                    Type: Expression(Integer)
--E

———-

2 [?]:14.637 ∫ csch^{2} ax dx

∫ csch^{2} ax = −\frac{coth ax}{a}

— * —

)clear all
--S 5 of 49
aa:=integrate(csch(a*x)^2,x)
--R
--R
--R                                     2
--R   (1)  - -------------------------------------------------------
--R                     2                                      2
--R          a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 6 of 49
bb:=-coth(a*x)/a
--R
--R          coth(a x)
--R   (2)  - ---------
--R              a
--R                                                    Type: Expression(Integer)
--E
--S 7 of 49      14:637 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                         2
--R       coth(a x)sinh(a x)  + 2cosh(a x)coth(a x)sinh(a x)
--R     +
--R                 2
--R       (cosh(a x)  - 1)coth(a x) - 2
--R  /
--R                2                                      2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a

--R                                                    Type: Expression(Integer)
--E

———-

3 [?]:14.638 ∫ csch^{3} ax dx

∫ csch^{3} ax = −\frac{csch ax coth ax}{2a} − \frac{1}{2a} ln tanh \frac{ax}{2}

— * —

)clear all
--S 8 of 49
aa:=integrate(csch(a*x)^3,x)
--R
--R
--R   (1)
--R                    4                      3              2              2
--R           sinh(a x)  + 4cosh(a x)sinh(a x)  + (6cosh(a x)  - 2)sinh(a x)
--R         +
--R                      3                                   4             2
--R           (4cosh(a x)  - 4cosh(a x))sinh(a x) + cosh(a x)  - 2cosh(a x)  + 1
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                      4                      3                2              2
--R           - sinh(a x)  - 4cosh(a x)sinh(a x)  + (- 6cosh(a x)  + 2)sinh(a x)
--R         +
--R                        3                                   4             2
--R           (- 4cosh(a x)  + 4cosh(a x))sinh(a x) - cosh(a x)  + 2cosh(a x)  - 1
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                   3                      2                2
--R       - 2sinh(a x)  - 6cosh(a x)sinh(a x)  + (- 6cosh(a x)  - 2)sinh(a x)
--R     +
--R                   3
--R       - 2cosh(a x)  - 2cosh(a x)
--R  /
--R                   4                        3                 2               2
--R       2a sinh(a x)  + 8a cosh(a x)sinh(a x)  + (12a cosh(a x)  - 4a)sinh(a x)
--R     +
--R                    3                                        4               2
--R       (8a cosh(a x)  - 8a cosh(a x))sinh(a x) + 2a cosh(a x)  - 4a cosh(a x)
--R     +
--R       2a
--R                                         Type: Union(Expression(Integer),...)

--E
--S 9 of 49
bb:=-(csch(a*x)*coth(a*x))/(2*a)-1/(2*a)*log(tanh((a*x)/2))
--R
--R                   a x
--R        - log(tanh(---)) - coth(a x)csch(a x)
--R                    2
--R   (2)  -------------------------------------
--R                          2a
--R                                                    Type: Expression(Integer)
--E
--S 10 of 49     14:638 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                    4                      3              2              2
--R           sinh(a x)  + 4cosh(a x)sinh(a x)  + (6cosh(a x)  - 2)sinh(a x)
--R         +
--R                      3                                   4             2
--R           (4cosh(a x)  - 4cosh(a x))sinh(a x) + cosh(a x)  - 2cosh(a x)  + 1
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                    4                      3              2              2
--R           sinh(a x)  + 4cosh(a x)sinh(a x)  + (6cosh(a x)  - 2)sinh(a x)
--R         +
--R                      3                                   4             2
--R           (4cosh(a x)  - 4cosh(a x))sinh(a x) + cosh(a x)  - 2cosh(a x)  + 1
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                      4                      3                2              2
--R           - sinh(a x)  - 4cosh(a x)sinh(a x)  + (- 6cosh(a x)  + 2)sinh(a x)
--R         +
--R                        3                                   4             2
--R           (- 4cosh(a x)  + 4cosh(a x))sinh(a x) - cosh(a x)  + 2cosh(a x)  - 1
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                                  4
--R       coth(a x)csch(a x)sinh(a x)
--R     +
--R                                                  3
--R       (4cosh(a x)coth(a x)csch(a x) - 2)sinh(a x)
--R     +
--R                   2                                              2

--R       ((6cosh(a x)  - 2)coth(a x)csch(a x) - 6cosh(a x))sinh(a x)
--R     +
--R                   3                                             2
--R       ((4cosh(a x)  - 4cosh(a x))coth(a x)csch(a x) - 6cosh(a x)  - 2)sinh(a x)
--R     +
--R               4             2                                    3
--R     (cosh(a x)  - 2cosh(a x)  + 1)coth(a x)csch(a x) - 2cosh(a x)  - 2cosh(a x)
--R  /
--R                   4                        3                 2               2
--R       2a sinh(a x)  + 8a cosh(a x)sinh(a x)  + (12a cosh(a x)  - 4a)sinh(a x)
--R     +
--R                    3                                        4               2
--R       (8a cosh(a x)  - 8a cosh(a x))sinh(a x) + 2a cosh(a x)  - 4a cosh(a x)
--R     +
--R       2a
--R                                                    Type: Expression(Integer)
--E

———-

4 [?]:14.639 ∫ csch^{n} ax coth ax dx

∫ csch ^{n}ax coth ax = −\frac{csch ^{n}ax}{na}

— * —

)clear all
--S 11 of 49
aa:=integrate(csch(a*x)^n*coth(a*x),x)
--R
--R
--R   (1)
--R                                 2sinh(a x) + 2cosh(a x)
--R       - sinh(n log(-------------------------------------------------))
--R                             2                                  2
--R                    sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R     +
--R                                 2sinh(a x) + 2cosh(a x)
--R       - cosh(n log(-------------------------------------------------))
--R                             2                                  2
--R                    sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R  /
--R     a n
--R                                         Type: Union(Expression(Integer),...)
--E
--S 12 of 49

bb:=-csch(a*x)^n/(n*a)
--R
--R                   n
--R          csch(a x)
--R   (2)  - ----------
--R              a n
--R                                                    Type: Expression(Integer)
--E
--S 13 of 49
cc:=aa-bb
--R
--R   (3)
--R                                 2sinh(a x) + 2cosh(a x)
--R       - sinh(n log(-------------------------------------------------))
--R                             2                                  2
--R                    sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R     +
--R                                 2sinh(a x) + 2cosh(a x)
--R       - cosh(n log(-------------------------------------------------))
--R                             2                                  2
--R                    sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R     +
--R                n
--R       csch(a x)
--R  /
--R     a n
--R                                                    Type: Expression(Integer)
--E
--S 14 of 49
cschrule:=rule(csch(x) == 1/sinh(x))
--R
--R                      1
--R   (4)  csch(x) == -------
--R                   sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 15 of 49
dd:=cschrule cc
--R
--R   (5)
--R                                 2sinh(a x) + 2cosh(a x)
--R       - sinh(n log(-------------------------------------------------))
--R                             2                                  2
--R                    sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R     +
--R                                 2sinh(a x) + 2cosh(a x)
--R       - cosh(n log(-------------------------------------------------))

--R                             2                                  2
--R                    sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R     +
--R            1     n
--R       (---------)
--R        sinh(a x)
--R  /
--R     a n
--R                                                    Type: Expression(Integer)
--E
--S 16 of 49
ee:=expandLog dd
--R
--R   (6)
--R       sinh
--R                           2                                  2
--R            n log(sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1)
--R          +
--R            - n log(sinh(a x) + cosh(a x)) - n log(2)
--R     +
--R       -
--R          cosh
--R                              2                                  2
--R               n log(sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1)
--R             +
--R               - n log(sinh(a x) + cosh(a x)) - n log(2)
--R     +
--R            1     n
--R       (---------)
--R        sinh(a x)
--R  /
--R     a n
--R                                                    Type: Expression(Integer)
--E
--S 17 of 49
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (7)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 18 of 49
ff:=sinhsqrrule ee
--R
--R   (8)
--R       sinh

--R                                                               2
--R                  4cosh(a x)sinh(a x) + cosh(2a x) + 2cosh(a x)  - 3
--R            n log(--------------------------------------------------)
--R                                           2
--R          +
--R            - n log(sinh(a x) + cosh(a x)) - n log(2)
--R     +
--R       -
--R          cosh
--R                                                                  2
--R                     4cosh(a x)sinh(a x) + cosh(2a x) + 2cosh(a x)  - 3
--R               n log(--------------------------------------------------)
--R                                              2
--R             +
--R               - n log(sinh(a x) + cosh(a x)) - n log(2)
--R     +
--R            1     n
--R       (---------)
--R        sinh(a x)
--R  /
--R     a n
--R                                                    Type: Expression(Integer)
--E
--S 19 of 49
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (9)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 20 of 49
gg:=coshsqrrule ff
--R
--R   (10)
--R       sinh
--R            n log(2cosh(a x)sinh(a x) + cosh(2a x) - 1)
--R          +
--R            - n log(sinh(a x) + cosh(a x)) - n log(2)
--R     +
--R       -
--R          cosh
--R               n log(2cosh(a x)sinh(a x) + cosh(2a x) - 1)
--R             +
--R               - n log(sinh(a x) + cosh(a x)) - n log(2)
--R     +
--R            1     n
--R       (---------)

--R        sinh(a x)
--R  /
--R     a n
--R                                                    Type: Expression(Integer)
--E
--S 21 of 49
sinhcoshrule:=rule(sinh(x)*cosh(y) == 1/2*(sinh(x+y)+sinh(x-y)))
--R
--I                              %O sinh(y + x) - %O sinh(y - x)
--I   (11)  %O cosh(y)sinh(x) == -------------------------------
--R                                             2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 22 of 49
hh:=sinhcoshrule gg
--R
--R   (12)
--R       sinh
--R            n log(sinh(2a x) + cosh(2a x) - 1) - n log(sinh(a x) + cosh(a x))
--R          +
--R            - n log(2)
--R     +
--R       -
--R          cosh
--R               n log(sinh(2a x) + cosh(2a x) - 1) - n log(sinh(a x) + cosh(a x))
--R             +
--R               - n log(2)
--R     +
--R            1     n
--R       (---------)
--R        sinh(a x)
--R  /
--R     a n
--R                                                    Type: Expression(Integer)
--E
--S 23 of 49     14:639 Schaums and Axiom agree
ii:=complexNormalize hh
--R
--R   (13)  0
--R                                                    Type: Expression(Integer)
--E

———-

5 [?]:14.640 ∫ \frac{dx}{csch ax} dx

∫ \frac{1}{csch ax} = \frac{1}{a}cosh ax

— * —

)clear all
--S 24 of 49
aa:=integrate(1/csch(a*x),x)
--R
--R
--R        cosh(a x)
--R   (1)  ---------
--R            a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 25 of 49
bb:=1/a*cosh(a*x)
--R
--R        cosh(a x)
--R   (2)  ---------
--R            a
--R                                                    Type: Expression(Integer)
--E
--S 26 of 49     14:640 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

6 [?]:14.641 ∫ x csch ax dx

∫ x csch ax = \frac{1}{a^{2}} \{ax − \frac{(ax)^{3}}{18} + \frac{7(ax)^{5}}{1800} + ⋯ + \frac{2(−1)^{n}(2^{2n−1} − 1)B_{n}(ax)^{2n+1}}{(2n + 1)!} + ⋯\}

— * —

)clear all
--S 27 of 49     14:641 Axiom cannot compute this integral

aa:=integrate(x*csch(a*x),x)
--R
--R
--R           x
--R         ++
--I   (1)   |   %O csch(%O a)d%O
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E

———-

7 [?]:14.642 ∫ x csch^{2} ax dx

∫ x csch^{2} ax = −\frac{x coth ax}{a} + \frac{1}{a^{2}} ln sinh ax

— * —

)clear all
--S 28 of 49
aa:=integrate(x*csch(a*x)^2,x)
--R
--R
--R   (1)
--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1)
--R      *
--R                     2sinh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                       2                                           2
--R       - 2a x sinh(a x)  - 4a x cosh(a x)sinh(a x) - 2a x cosh(a x)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 29 of 49
bb:=-(x*coth(a*x))/a+1/a^2*log(sinh(a*x))
--R
--R        log(sinh(a x)) - a x coth(a x)
--R   (2)  ------------------------------
--R                       2
--R                      a
--R                                                    Type: Expression(Integer)

--E
--S 30 of 49
cc:=aa-bb
--R
--R   (3)
--R                   2                                  2
--R       (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  + 1)log(sinh(a x))
--R     +
--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1)
--R      *
--R                     2sinh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                                      2
--R       (a x coth(a x) - 2a x)sinh(a x)
--R     +
--R       (2a x cosh(a x)coth(a x) - 4a x cosh(a x))sinh(a x)
--R     +
--R                     2                                 2
--R       (a x cosh(a x)  - a x)coth(a x) - 2a x cosh(a x)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                                    Type: Expression(Integer)
--E
--S 31 of 49
dd:=expandLog cc
--R
--R   (4)
--R                     2                                  2
--R         (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  + 1)
--R      *
--R         log(sinh(a x) - cosh(a x))
--R     +
--R                                                 2
--R       (a x coth(a x) + log(- 2) - 2a x)sinh(a x)
--R     +
--R       (2a x cosh(a x)coth(a x) + (2log(- 2) - 4a x)cosh(a x))sinh(a x)
--R     +
--R                     2                                             2
--R       (a x cosh(a x)  - a x)coth(a x) + (log(- 2) - 2a x)cosh(a x)  - log(- 2)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                                    Type: Expression(Integer)
--E

--S 32 of 49
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (5)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 33 of 49
ee:=sinhsqrrule dd
--R
--R   (6)
--R                                                         2
--R         (- 4cosh(a x)sinh(a x) - cosh(2a x) - 2cosh(a x)  + 3)
--R      *
--R         log(sinh(a x) - cosh(a x))
--R     +
--R       (4a x cosh(a x)coth(a x) + (4log(- 2) - 8a x)cosh(a x))sinh(a x)
--R     +
--R                                       2
--R       (a x cosh(2a x) + 2a x cosh(a x)  - 3a x)coth(a x)
--R     +
--R                                                                2
--R       (log(- 2) - 2a x)cosh(2a x) + (2log(- 2) - 4a x)cosh(a x)  - 3log(- 2)
--R     +
--R       2a x
--R  /
--R       2                      2               2         2     2
--R     4a cosh(a x)sinh(a x) + a cosh(2a x) + 2a cosh(a x)  - 3a
--R                                                    Type: Expression(Integer)
--E
--S 34 of 49
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (7)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 35 of 49
ff:=coshsqrrule ee
--R
--R   (8)
--R       (- 2cosh(a x)sinh(a x) - cosh(2a x) + 1)log(sinh(a x) - cosh(a x))
--R     +
--R       (2a x cosh(a x)coth(a x) + (2log(- 2) - 4a x)cosh(a x))sinh(a x)

--R     +
--R       (a x cosh(2a x) - a x)coth(a x) + (log(- 2) - 2a x)cosh(2a x) - log(- 2)
--R  /
--R       2                      2              2
--R     2a cosh(a x)sinh(a x) + a cosh(2a x) - a
--R                                                    Type: Expression(Integer)
--E
--S 36 of 49
sinhcoshrule:=rule(sinh(x)*cosh(y) == 1/2*(sinh(x+y)+sinh(x-y)))
--R
--I                             %P sinh(y + x) - %P sinh(y - x)
--I   (9)  %P cosh(y)sinh(x) == -------------------------------
--R                                            2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 37 of 49
gg:=sinhcoshrule ff
--R
--R   (10)
--R       (- sinh(2a x) - cosh(2a x) + 1)log(sinh(a x) - cosh(a x))
--R     +
--R       (a x coth(a x) + log(- 2) - 2a x)sinh(2a x)
--R     +
--R       (a x cosh(2a x) - a x)coth(a x) + (log(- 2) - 2a x)cosh(2a x) - log(- 2)
--R  /
--R      2              2              2
--R     a sinh(2a x) + a cosh(2a x) - a
--R                                                    Type: Expression(Integer)
--E
--S 38 of 49     14:642 Schaums and Axiom differ by a constant
hh:=complexNormalize gg
--R
--R         - log(- 1) + log(- 2)
--R   (11)  ---------------------
--R                    2
--R                   a
--R                                                    Type: Expression(Integer)
--E

———-

8 [?]:14.643 ∫ \frac{csch ax}{x} dx

∫ \frac{csch ax}{x} = −\frac{1}{ax} − \frac{ax}{6} + \frac{7(ax)^{3}}{1080} + ⋯ \frac{(−1)^{n}2(2^{2n−1} − 1)B_{n}(ax)^{2n−1}}{(2n − 1)(2n)!} + ⋯

— * —

)clear all
--S 39 of 49     14:643 Axiom cannot compute this integral
aa:=integrate(csch(a*x)/x,x)
--R
--R
--R           x
--I         ++  csch(%O a)
--I   (1)   |   ---------- d%O
--I        ++       %O
--R                                         Type: Union(Expression(Integer),...)
--E

———-

9 [?]:14.644 ∫ \frac{dx}{q + p csch ax} dx

∫ \frac{1}{q + p csch ax} = \frac{x}{q} − \frac{p}{q} ∫ \frac{dx}{p + q sinh ax}

— * —

)clear all
--S 40 of 49
aa:=integrate(1/(q+p*csch(a*x)),x)
--R
--R
--R   (1)
--R         p
--R      *
--R         log
--R                     2         2      2                              2         2
--R                    q sinh(a x)  + (2q cosh(a x) + 2p q)sinh(a x) + q cosh(a x)
--R                  +
--R                                      2     2
--R                    2p q cosh(a x) + q  + 2p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                   3     2                 3     2                  2     3
--R                (2q  + 2p q)sinh(a x) + (2q  + 2p q)cosh(a x) + 2p q  + 2p
--R           /

--R                           2                                             2
--R                q sinh(a x)  + (2q cosh(a x) + 2p)sinh(a x) + q cosh(a x)
--R              +
--R                2p cosh(a x) - q
--R     +
--R           +-------+
--R           | 2    2
--R       a x\|q  + p
--R  /
--R         +-------+
--R         | 2    2
--R     a q\|q  + p
--R                                         Type: Union(Expression(Integer),...)
--E
--S 41 of 49
t1:=integrate(1/(p+q*sinh(a*x)),x)
--R
--R   (2)
--R     log
--R                 2         2      2                              2         2
--R                q sinh(a x)  + (2q cosh(a x) + 2p q)sinh(a x) + q cosh(a x)
--R              +
--R                                  2     2
--R                2p q cosh(a x) + q  + 2p
--R           *
--R               +-------+
--R               | 2    2
--R              \|q  + p
--R          +
--R                 3     2                   3     2                  2     3
--R            (- 2q  - 2p q)sinh(a x) + (- 2q  - 2p q)cosh(a x) - 2p q  - 2p
--R       /
--R                       2                                             2
--R            q sinh(a x)  + (2q cosh(a x) + 2p)sinh(a x) + q cosh(a x)
--R          +
--R            2p cosh(a x) - q
--R  /
--R       +-------+
--R       | 2    2
--R     a\|q  + p
--R                                         Type: Union(Expression(Integer),...)
--E
--S 42 of 49
bb:=x/q-p/q*t1
--R
--R   (3)
--R       -
--R            p

--R         *
--R            log
--R                        2         2      2
--R                       q sinh(a x)  + (2q cosh(a x) + 2p q)sinh(a x)
--R                     +
--R                        2         2                     2     2
--R                       q cosh(a x)  + 2p q cosh(a x) + q  + 2p
--R                  *
--R                      +-------+
--R                      | 2    2
--R                     \|q  + p
--R                 +
--R                      3     2                   3     2                  2     3
--R                 (- 2q  - 2p q)sinh(a x) + (- 2q  - 2p q)cosh(a x) - 2p q  - 2p
--R              /
--R                              2                                             2
--R                   q sinh(a x)  + (2q cosh(a x) + 2p)sinh(a x) + q cosh(a x)
--R                 +
--R                   2p cosh(a x) - q
--R     +
--R           +-------+
--R           | 2    2
--R       a x\|q  + p
--R  /
--R         +-------+
--R         | 2    2
--R     a q\|q  + p
--R                                                    Type: Expression(Integer)
--E
--S 43 of 49
cc:=aa-bb
--R
--R   (4)
--R         p
--R      *
--R         log
--R                     2         2      2                              2         2
--R                    q sinh(a x)  + (2q cosh(a x) + 2p q)sinh(a x) + q cosh(a x)
--R                  +
--R                                      2     2
--R                    2p q cosh(a x) + q  + 2p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                   3     2                 3     2                  2     3
--R                (2q  + 2p q)sinh(a x) + (2q  + 2p q)cosh(a x) + 2p q  + 2p
--R           /

--R                           2                                             2
--R                q sinh(a x)  + (2q cosh(a x) + 2p)sinh(a x) + q cosh(a x)
--R              +
--R                2p cosh(a x) - q
--R     +
--R         p
--R      *
--R         log
--R                     2         2      2                              2         2
--R                    q sinh(a x)  + (2q cosh(a x) + 2p q)sinh(a x) + q cosh(a x)
--R                  +
--R                                      2     2
--R                    2p q cosh(a x) + q  + 2p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                     3     2                   3     2                  2     3
--R                (- 2q  - 2p q)sinh(a x) + (- 2q  - 2p q)cosh(a x) - 2p q  - 2p
--R           /
--R                           2                                             2
--R                q sinh(a x)  + (2q cosh(a x) + 2p)sinh(a x) + q cosh(a x)
--R              +
--R                2p cosh(a x) - q
--R  /
--R         +-------+
--R         | 2    2
--R     a q\|q  + p
--R                                                    Type: Expression(Integer)
--E
--S 44 of 49
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (5)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 45 of 49
dd:=sinhsqrrule cc
--R
--R   (6)
--R         p
--R      *
--R         log
--R                       2                              2
--R                    (4q cosh(a x) + 4p q)sinh(a x) + q cosh(2a x)

--R                  +
--R                      2         2                     2     2
--R                    2q cosh(a x)  + 4p q cosh(a x) + q  + 4p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                   3     2                 3     2                  2     3
--R                (4q  + 4p q)sinh(a x) + (4q  + 4p q)cosh(a x) + 4p q  + 4p
--R           /
--R                                                                          2
--R                (4q cosh(a x) + 4p)sinh(a x) + q cosh(2a x) + 2q cosh(a x)
--R              +
--R                4p cosh(a x) - 3q
--R     +
--R         p
--R      *
--R         log
--R                       2                              2
--R                    (4q cosh(a x) + 4p q)sinh(a x) + q cosh(2a x)
--R                  +
--R                      2         2                     2     2
--R                    2q cosh(a x)  + 4p q cosh(a x) + q  + 4p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                     3     2                   3     2                  2     3
--R                (- 4q  - 4p q)sinh(a x) + (- 4q  - 4p q)cosh(a x) - 4p q  - 4p
--R           /
--R                                                                          2
--R                (4q cosh(a x) + 4p)sinh(a x) + q cosh(2a x) + 2q cosh(a x)
--R              +
--R                4p cosh(a x) - 3q
--R  /
--R         +-------+
--R         | 2    2
--R     a q\|q  + p
--R                                                    Type: Expression(Integer)
--E
--S 46 of 49
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (7)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))

--E
--S 47 of 49
ee:=coshsqrrule dd
--R
--R   (8)
--R         p
--R      *
--R         log
--R                       2                              2
--R                    (2q cosh(a x) + 2p q)sinh(a x) + q cosh(2a x)
--R                  +
--R                                      2     2
--R                    2p q cosh(a x) + q  + 2p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                   3     2                 3     2                  2     3
--R                (2q  + 2p q)sinh(a x) + (2q  + 2p q)cosh(a x) + 2p q  + 2p
--R           /
--R              (2q cosh(a x) + 2p)sinh(a x) + q cosh(2a x) + 2p cosh(a x) - q
--R     +
--R         p
--R      *
--R         log
--R                       2                              2
--R                    (2q cosh(a x) + 2p q)sinh(a x) + q cosh(2a x)
--R                  +
--R                                      2     2
--R                    2p q cosh(a x) + q  + 2p
--R               *
--R                   +-------+
--R                   | 2    2
--R                  \|q  + p
--R              +
--R                     3     2                   3     2                  2     3
--R                (- 2q  - 2p q)sinh(a x) + (- 2q  - 2p q)cosh(a x) - 2p q  - 2p
--R           /
--R              (2q cosh(a x) + 2p)sinh(a x) + q cosh(2a x) + 2p cosh(a x) - q
--R  /
--R         +-------+
--R         | 2    2
--R     a q\|q  + p
--R                                                    Type: Expression(Integer)
--E
--S 48 of 49     14:644 Schaums and Axiom differ by a constant
ff:=complexNormalize ee

--R
--R               4    2 2
--R        p log(q  + p q )
--R   (9)  ----------------
--R              +-------+
--R              | 2    2
--R          a q\|q  + p
--R                                                    Type: Expression(Integer)
--E

———-

10 [?]:14.645 ∫ csch^{n} ax dx

∫ csch^{n} ax = \frac{−csch^{n−2} ax coth ax}{a(n − 1)} − \frac{n − 2}{n − 1} ∫ csch^{n−2} ax

— * —

)clear all
--S 49 of 49     14:645 Axiom cannot compute this integral
aa:=integrate(csch(a*x)^n,x)
--R
--R
--R           x
--R         ++            n
--I   (1)   |   csch(%O a) d%O
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E
)spool
)lisp (bye)

———-

References

[1] Spiegel, Murray R. Mathematical Handbook of Formulas and Tables Schaum’s Outline Series McGraw-Hill 1968 pp91-92